Insert two geometric means between 4 and 500.
a1=4
a2=
a3=
a=500
1
Expert's answer
2012-12-28T04:42:26-0500
let q is parameter
then a1=4, a2=4q, a3=4q^2, a4=4q^3=500
we have that 4q^3=500 q^3=125 q=5
so a2=4*5=20, a3=4*25=100